OpenStudy (anonymous):

use the quadratic formula to solve 2y=4y^2-6

OpenStudy (anonymous):

first put it in the form ax^2 + bx + c = 0 i.e 4y^2 -2y -6 = 0 here a = 4, b = -2 and c = -6

OpenStudy (anonymous):

quadratic formula is \[- b \pm \sqrt{(b ^{2}}-4ac)/2a\]

OpenStudy (anonymous):

so - (-2) +/- root[(-2)^2 - 4*(4)*(-6)] 2 +/- root[4 + 96] x = ------------------------------- = -------------------- 2 * 4 8 2 +/- root[100] 2 +/- 10 x = --------------- = --------- 8 8

OpenStudy (anonymous):

so 2 + 10 2 - 10 12 -8 x = -------- or -------- = ----- or -------- 8 8 8 8 so x = 3/2 or x = -1

OpenStudy (anonymous):

Hey friend, pls click GOOD ANSWER button !!!! ( :{)

OpenStudy (anonymous):

\[2y=4y^2-6\] \[4y^2-2y-6=0\] \[2y^2-y-3=0\] \[y=\frac{1\pm\sqrt{1-4\times 2\times -3}}{2\times 2}\] \[y=\frac{1\pm\sqrt{25}}{4}\] \[y=\frac{1+5}{4}=\frac{6}{4}=\frac{3}{2}\] or \[y=\frac{1-5}{4}=\frac{-4}{4}=-1\]